Mattola Company is giving each of its employees a holiday bonus of $900 on December 14 (a non payday). The company wants each employee's check to be $900. The supplemental tax percent is used. Nobody has capped for OASDI prior to the bonus check.

a. What will be the gross amount of each bonus if each employee pays a state income tax of 2.8% (besides the other payroll taxes)? You may need to add one penny to the gross so that net bonus exactly equals $900.

b. What would the net amount of each bonus check be if the company did not gross-up the bonus?
Answer #1

a ) Calculation of gross amount of each bonus

state income tax = 2.8 %

FICA social security tax = 6.20 %

FICA medicare tax = 1.45 %

gross amount of each bonus = 900 / ( 100 % - 25 % - 6.20 % - 1.45 % - 2.8 % )

= 900/64.55%

= 1394.27

b ) Calculation of net amount of each bonus

= 500 * ( 100 % - 25 % - 6.20 % - 1.45 % - 2.80 % )

= 900* 64.55 %

= 580.95